OpenStudy (anonymous):

Range of score help You are studying for your final exam of the semester. Up to this point, you received 3 exam scores of 83%, 85%, and 84%. To receive a grade of B in the class, you must have an average exam score between 80% and 89% for all 4 exams including the final. Find the widest range of scores that you can get on the final exam in order to receive a grade of B for the class. Use interval notation. a 71,72 b 62,100 c 68,100

OpenStudy (anonymous):

To find the mean/average score, you need to add all of the scores and divide by the number of scores you have. If you make the final exam (the unknown), s, then you get the following formula for the final average: (83+85+84+s)/4=f, where f=final score Simplified (add the already known scores together), you get, (252+s)/4=f The lowest final score is 80 and the highest is 89, so you just have to find the value of s when f=80 and f=89 as follows: (252+s)/4=80 and (252+s)/4=89 further, you try to solve, if u didnt get, then ask.